Co(ll) was used as an internal standard to analyze a sample of Ti(IV) with atomic absorption spectroscopy (AAS). A standard mixture containing 1.32 ug Co/mL and 1.54 ug TimL measured by AAS produced a signal-to-signal ratio of 1.78 Ti: 1.00 Co A mixture was prepared by combining 5.00 mL of a Ti(IV) solution of unknown concentration with 3.50 mL of a 11.7 ug/mL solution of Co (ll). The absorbance of the mixture at the Ti(IV) wavelength was 0.149 and the absorbance at the Co (ll) wavelength was 0.202. Determine the concentration, in moles per liter, of Ti(IV) in the original unknown solution. Number M Ti (IV)
Answer #1

STANDARD SOLUTION IS PREPARED WITH Co AND Ti WITH

cO :Ti :: 1.00 : 1.78

THEN Ti / cO is 1.78

AND THE CONCENTRATIONS USED ARE

Ti IS 1.32  µg/ml AND Co IS 1.54  µg/ml.

IN THE PRESENT EXPERIMENT, 3.5/ml IS MIXED WITH 5.00 ml OF UNKNOWN CONCENTRATION OFTi.

THE CONCENTRATION OF THE GIVEN Ti IN UNKNOWN SOLUTION IS

Ti IS 1.78 * 1.149 / 0.202 = 10.12

THE FORMUALA CAN BE ARRIVED BY CALCULATING 5.00 * 10.12 / 3.50 = 14.57

HENCE HE CONCENTRATION OF i IS 14.57 * 0.149 / 0.202 = 10.74  µg / ml

µ